OpenStudy (anonymous):

what are (x-4)^2/(36)-(y-2)^2/(9)=1 asymptotes?

OpenStudy (anonymous):

hyperbola?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[\huge \frac{(x-4)^2}{36}-\frac{(y-2)^2}{9}=1 \]

OpenStudy (anonymous):

according to your equation above, what is the center here?

OpenStudy (anonymous):

(4,2)

OpenStudy (anonymous):

ok.. so we KNOW that both asymptotes go through this center... all we need are the slopes... read here, page 3, about the slopes of hyperbolas: http://tutorial.math.lamar.edu/pdf/Algebra_Cheat_Sheet.pdf

OpenStudy (anonymous):

then see if you can come up with the equation of the asymptotes of your hyperbola....

OpenStudy (anonymous):

since this hyperbola open left and right so equation of asymptotic is \[\Large y=\pm(\frac{b}{a}) x\] where b=3 a=6 so \[\Large y=\pm(\frac{3}{6}) x=\pm 0.5x\]

OpenStudy (anonymous):

@SantanaG did you get it ?

OpenStudy (anonymous):

yeah thanks @sami-21

OpenStudy (anonymous):

Here are the equations of the asymptotes \[ y-2 =\pm \frac 1 2 (x-4) \]